OpenStudy (trojanpoem):

lim_{x rightarrow 2}\frac{ 3^{x} - 9}{ x - 2}

OpenStudy (gorv):

\[\lim_{x \rightarrow 2}\frac{ 3^x-9 }{ x-2 }\]

OpenStudy (gorv):

when we put x=2 in it it give \[\frac{ 0 }{ 0 }\]

OpenStudy (gorv):

so we can use l hopital rule here

OpenStudy (gorv):

which means differentiate numerator and denominator seprately

OpenStudy (gorv):

\[\lim_{x \rightarrow 2}\frac{ \frac{ d(3^x-9) }{ dx } }{ \frac{ d(x-2) }{ dx } }\]

OpenStudy (gorv):

can u differentiate it now??

OpenStudy (trojanpoem):

Can you solve it to me using logs ?